Explain the differences between solving these two equations:
• log3(x - 1) = 4 and log3(x - 1) = log34

Answer by ewatrrr(24785)


Hi,
log3(x - 1) = 4
Using the following Property of Logarithms

3^4 = (x-1)
81 = x-1
82 = x
log3(x - 1) = log34 logarithm bases are equal in this example, therefore:
(x-1) = 4
x = 5